# Cliqueformer: Model-Based Optimization with Structured Transformers

### v1

@@ 3.2 Architecture
The goal of this subsection is to derive an MBO model that can simultaneously learn the target function as well as its structure, and thus be readily applied to MBO. First, we would like the model to decompose its prediction into a sum of models defined over small subsets of the input variables in the manner of Equation (3). As discussed in the previous subsection, efforts to discover such a structure are impractical since there exists a plethora of reparameterizations of the data and their corresponding FGMs of varied decomposabilities. Thus, instead, we propose that the FGM be defined first, and a representation of the data be learned to align with the chosen structure.
@@
Accomplishing this desideratum is simple. A model that we train should, after some transformations, split the input’s representations, denoted as ${\mathbf{z}}$, into partially-overlapping cliques, $({\mathbf{z}}_{C_{1}},\dots,{\mathbf{z}}_{C_{N_{\text{clique}}}})$, and process them independently, followed by a summation. To implement this, we specify how many cliques we want to decompose the function into, $N_{\text{clique}}$, as well as their dimensionality, $d_{\text{clique}}$, and the size of their knots—dimensions at which two consecutive cliques overlap, $d_{\text{knot}}=|{\mathbf{z}}_{C_{i}}\cap{\mathbf{z}}_{C_{i+1}}|$. Then, we pass each clique through an MLP network that is also equipped with a trigonometric clique embedding, similar to vaswani2017attention to express a different function for each clique,
$f_{\theta}({\mathbf{z}})=\frac{1}{N_{\text{clique}}}\sum_{i=1}^{N_{\text{clique}}}f_{\theta}({\mathbf{z}}_{C_{i}},{\bm{c}}_{i})$,  where  $\begin{bmatrix}{\bm{c}}_{i,2j}&{\bm{c}}_{i,2j+1}\end{bmatrix}=\begin{bmatrix}\sin(i\cdot\omega_{j})&\cos(i\cdot\omega_{j})\end{bmatrix}$,
and $\omega_{j}=10^{-8j/d_{\text{model}}}$. Here, we use the arithmetic mean over cliques rather than summation because, while being functionally equivalent, it provides more stability when $N_{\text{clique}}\rightarrow\infty$. Pre-defining the cliques over the representations allows us to avoid the problem of discovering arbitrarily dense graphs, as explained in Subsection 3.1. This architectural choice implies that the regret from Theorem 1 will depend on the coverage of such representations’ cliques, $\max_{i\in[N_{\text{clique}}]}1/e_{\theta}({\mathbf{z}}_{C_{i}})$, where $e_{\theta}({\mathbf{z}})=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}[e_{\theta}({\mathbf{z}}|{\mathbf{x}})]$ is the marginal distribution of the representations learned by an encoder $e_{\theta}$. This term can become dangerously large if individual distributions $e_{\theta}({\mathbf{z}}_{C})$ put disproportionally more density to some regions of the latent space than to others, Thus, to prevent that, we propose to train the latent space so that the distribution of cliques attain wide coverage.
@@
The model should learn representations whose cliques have dsirtibutions featured by wide coverage.
@@
To meet this requirement, we leverage tools from representation learning, but in a novel way. Namely, we put a variational bottleneck (kingma2013auto; higgins2016beta; alemi2016deep, VIB) on individual cliques of our representations that brings their distribution closer to a prior with wide coverage, which we choose to be the standard-normal prior. To implement it, when computing the loss for a single example, we sample a single clique to compute the VIB for at random, as opposed to computing it for the joint latent variable like in the classical VIB,
$\text{VIB}({\mathbf{x}},\theta)=\mathbb{E}_{i\sim U[N_{\text{clique}}]}\big{[}\text{KL}\big{(}e_{\theta}({\mathbf{z}}_{C_{i}}|{\mathbf{x}}),p_{C_{i}}({\mathbf{z}}_{C_{i}})\big{)}\big{]},$ (4)
@@
where $e_{\theta}({\mathbf{z}}_{C}|{\mathbf{x}})$ is the density of clique $C$ produced by our learnable encoder $e_{\theta}({\mathbf{z}}|{\mathbf{x}})$, and $p_{C}({\mathbf{z}}_{C})$ is the density of ${\mathbf{z}}_{C}$ under the standard-normal distrituion. Note that it is not equivalent to the classical, down-weighted VIB in expectation either since our cliques overlap, meaning that knots contribute to the VIB more often than regular dimensions. Together with the model $f_{\theta}$ and the encoder $e_{\theta}$, we train a decoder $d_{\theta}({\mathbf{x}}|{\mathbf{z}})$ that reconstructs the designs from the latent variables. Putting it all together, the training objective of our model is a VIB-style likelihood objective with a regression term,
@@
$L_{\text{clique}}(\theta)=\mathbb{E}_{({\mathbf{x}},{\textnormal{y}})\sim\mathcal{D},{\mathbf{z}}\sim e_{\theta}(\cdot|{\mathbf{z}}),i\sim U[N_{\text{clique}}]}\Big{[}\text{VIB}({\mathbf{x}},\theta)-\log d_{\theta}({\mathbf{x}}|{\mathbf{z}})+\tau\cdot\big{(}{\textnormal{y}}-f_{\theta}({\mathbf{z}})\big{)}^{2}\Big{]},$ (5)
where $\tau$ is a positive coefficient that we set to $10$ in our experiments.
@@
Figure 4: Illustration of information flow in Cliqueformer’s training. Data are shown in navy, learnable variables in blue, neural modules in pink, and loss functions in green. The input ${\mathbf{x}}$ is passed to a transformer encoder to compute representation ${\mathbf{z}}$ which is decomposed into cliques with small overlapping knots (highlighted in colors on the figure). The representation goes to the parallel MLPs whose outputs, added together, predict target y. The representation ${\mathbf{z}}$ is also fed to a transformer decoder that tries to recover the original input ${\mathbf{x}}$. Additionally, the representation goes through an information bottleneck from Equation (4) during training.
@@
Since in our neural network we impose the FGM decomposition of the predictive module in the latent space, we need to endow our model with high expressivity to learn representations that meet such demands. Thus, we model our encoder $e_{\theta}$ and decoder $d_{\theta}$ with transformer networks (vaswani2017attention). To leverage such an architecture, the encoder begins by mapping the input vector ${\mathbf{x}}\in\mathbb{R}^{d}$ into $d$ vectors of dimensionality $d_{\text{model}}$, which it then processes as if they were a sequence of token embeddings. After a series of transformer blocks, the network then maps the sequence into a normal distribution over representations, $e_{\theta}({\mathbf{z}}|{\mathbf{x}})$. A vector ${\mathbf{z}}$ can then be sampled from that distribution and arranged into $N_{\text{clique}}$ cliques with dimensionality $d_{\text{clique}}$ and knot size of $d_{\text{knot}}$. The new sequence can then be fed into the predictive model $f_{\theta}$ and to the decoder $d_{\theta}$, which is a transformer too. For illustration of the information flow in Cliqueformer’s training consult Figure 4.
@@
3:     Take a gradient step on the parameter $\theta$ with respect to $L_{\text{clique}}(\theta)$ from Equation (5).
@@
5:  Sample $B$ examples ${\mathbf{x}}^{(i_{b})}\sim\mathcal{D}$, $b\in[B]$, from the dataset.
6:  Encode the examples with the encoder ${\mathbf{z}}^{(i_{b})}\sim e_{\theta}({\mathbf{z}}|{\mathbf{x}}^{(i_{b})})$, for $b\in[B]$.
@@
8:     Decay the representation ${\mathbf{z}}$ of the design, ${\mathbf{z}}^{(i_{b})}\leftarrow(1-\lambda){\mathbf{z}}^{(i_{b})},\ \forall b\in[B]$.
@@ 3.3 Optimizing Designs With Cliqueformer
Once Cliqueformer is trained, we use it to optimize new designs. Typically, MBO methods initialize this step at a sample of designs $({\mathbf{x}}^{i_{b}})_{b=1}^{B}$ drawn from the dataset (trabucco2021conservative). Since in our algorithm the optimization takes place in the latent space $\mathcal{Z}$, we perform this step by encoding the sample of designs with Cliqueformer’s encoder, ${\mathbf{z}}^{i_{b}}\sim e_{\theta}({\mathbf{z}}|{\mathbf{x}}^{i_{b}})$. We then optimize the representation ${\mathbf{z}}^{i_{b}}$ of design ${\mathbf{x}}^{i_{b}}$ to maximize our model’s value,
$L_{\text{mbo}}\big{(}({\mathbf{z}}^{i_{b}})_{b=1}^{B}\big{)}=\frac{1}{B}\sum_{b=1}^{B}f_{\theta}({\mathbf{z}}^{i_{b}}),$ (6)
at the same time minding the enumerator of the regret bound from Theorem 1. That is, we don’t want the optimizer to explore regions under which the marginal densities $e_{\theta}({\mathbf{z}}_{C})=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}[e_{\theta}({\mathbf{z}}_{C}|{\mathbf{x}})]$ are small. Fortunately, since the encoder was trained with standard-normal prior on the cliques, $p({\mathbf{z}}_{C})=N(0_{d_{\text{clique}}},I_{d_{\text{clique}}})$, we know that values of ${\mathbf{z}}$ closer to the origin have unilaterally higher marginals. This simple property of standard-normal distribution allows us to confine the optimizer’s exploration to designs with in-distribution cliques by exponentially decaying the design at every optimization step. Thus, we use AdamW as our optimizer (loshchilov2017fixing). We provide the pseudocode of the whole procedure of designign with Cliqueformer in Algorithm 1.
@@ 4 Related Work
The idea of using machine learning models in optimization problems has existed for a long time, and has been mainly cultivated in the literature on Bayesian optimization (williams2006gaussian; brochu2010tutorial; snoek2012practical, BO). The BO paradigm relies on two core assumptions: availability of data of examples paired with their target function values, as well as access to an oracle that allows a learning algorithm to query values of proposed examples. Thus, similarly to reinforcement learning, the challenge of BO is to balance exploitation and exploration of the black-box function modeled by a Gaussian process.
@@
Recently, to help BO tackle very high-dimensional problems, techniques of decomposing the target function have become more popular (kandasamy2015high; rolland2018high). Most commonly, these methods decompose the target function into functions defined on the input’s partitions. While such models are likely to deviate far from the functions’ ground-truth structure, one can derive theoretical guarantees for a range of such decompositions under the BO’s query budget assumptions (ziomek2023random). However, these results do not apply to our setting of offline MBO, where no additional queries are available and the immediate reliability on the model is essential. Furthermore, instead of partitioning the input, we decompose our prediction over a latent variable that is learned by a transformer, enabling the model to acquire an expressive structure over which the decomposition is valid.
@@
Offline model-based optimization (MBO) has been recently attracting attention of researchers and practitioners from domains where BO assumptions cannot be easily met, offering an attractive premise of producing solutions directly after training on static datasets, without the need for additional queries. One of the first tasks tackled by MBO was molecule design (gomez2018automatic), where a variational auto-encoder (kingma2013auto, VAE) was used to learn continuous representations of molecular data. In contrast to our work, however, this work does not study learning structural properties of the target function, but rather is a proof of concept of applying deep learning to molecular design. A data type-agnostic algorithm was introduced by brookes2019conditioning, dubbed Conditioning by Adaptive Sampling (CBaS), that iteratively refines its design proposals in response to predictions of a non-differentiable oracle. While one can use this refinement procedure in combination with trainable models by means of auto-focusing (fannjiang2020autofocused), this setting is different than ours since we assume the ability to model the black-box function with a neural network. Since such a model is differentiable, we can simply rely on automatic differentiation (paszke2019pytorch) to refine our designs. trabucco2021conservative introduced a neural network-based method, exactly for our setting, dubbed Conservative Objective Models (COMs), where a surrogate model is trained to both predict values of examples that can be found in the dataset, and penalize those that are not. COMs differs from our work fundamentally, since its contribution lies in the formulation of the conservative regularizer applied to arbitrary neural networks, while we focus on scalable model architectures that facilitate computational design. Another recent line of work proposes to tackle the design problem through means of generative modeling. BONET (mashkaria2023generative) and DDOM (pmlr-v202-krishnamoorthy23a) are examples of works that bring the most recent novelties of the field to address design tasks. BONET does so by training a transformer to generate sequences of designs that monotonically improve in their value, and DDOM by training a value-conditioned diffusion model. That is, these methods attempt to generate high-value designs through novel conditional generation mechanisms. Instead, we model MBO as a maximization problem, and propose a scalable model that acquires the structure of the black-box function through standard gradient-based learning. To this end, we bring powerful techniques from deep learning and generative modeling, like transformers (vaswani2017attention) and variational-information bottlenecks (kingma2013auto; alemi2016deep).
@@
The work of grudzien2024functional introduced the theoretical foundations of functional graphical models (FGMs), including Theorem 1. However, as we have shown in Theorem 2, their graph discovery heuristic for neural networks renders learning the black-box function’s structure an open problem. Our work addresses this issue by subsuming the graph discovery step in the architecture of our Cliqueformer that learns to abide by a pre-defined FGM. As such, the model learns the structure of the target function, as well as learns to predict its value, in synergy within an end-to-end training.

### v2

@@ 3.2 Architecture
The goal of this subsection is to derive an MBO model that can simultaneously learn the target function as well as its structure, and thus be readily applied to MBO. First, we would like the model to decompose its prediction into a sum of models defined over small subsets of the input variables in the manner of Equation (3). As discussed in the previous subsection, in general settings, while such a structure is not known a priori, a sufficiently expressive model can learn to follow a pre-defined structure. Thus, instead, we propose that the FGM be defined first, and a representation of the data be learned to align with the chosen structure.
@@
Accomplishing this desideratum is simple. A model that we train should, after some transformations, split the input’s representations, denoted as ${\mathbf{z}}$, into partially-overlapping cliques, $({\mathbf{z}}_{C_{1}},\dots,{\mathbf{z}}_{C_{N_{\text{clique}}}})$, and process them independently, followed by a summation. To implement this, we specify how many cliques we want to decompose the function into, $N_{\text{clique}}$, as well as their dimensionality, $d_{\text{clique}}$, and the size of their knots—dimensions at which two consecutive cliques overlap, $d_{\text{knot}}=|{\mathbf{z}}_{C_{i}}\cap{\mathbf{z}}_{C_{i+1}}|$. Then, we pass each clique through an MLP network that is also equipped with a trigonometric clique embedding, similar to Vaswani et al. (2017) to express a different function for each clique,
$f_{\theta}({\mathbf{z}})=\frac{1}{N_{\text{clique}}}\sum_{i=1}^{N_{\text{clique}}}f_{\theta}({\mathbf{z}}_{C_{i}},{\bm{c}}_{i})$,
where  ${\bm{c}}_{i,2j},{\bm{c}}_{i,2j+1}=\sin(i\cdot\omega_{j}),\cos(i\cdot\omega_{j})$
and $\omega_{j}=10^{-8j/d_{\text{model}}}$. Here, we use the arithmetic mean over cliques rather than summation because, while being functionally equivalent, it provides more stability when $N_{\text{clique}}\rightarrow\infty$. Pre-defining the cliques over the representations allows us to avoid the problem of discovering arbitrarily dense graphs, as explained in Subsection 3.1. This architectural choice implies that the regret from Theorem 1 will depend on the coverage of such representations’ cliques, $\max_{i\in[N_{\text{clique}}]}1/e_{\theta}({\mathbf{z}}_{C_{i}})$, where $e_{\theta}({\mathbf{z}})=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}[e_{\theta}({\mathbf{z}}|{\mathbf{x}})]$ is the marginal distribution of the representations learned by an encoder $e_{\theta}$. This term can become dangerously large if individual distributions $e_{\theta}({\mathbf{z}}_{C})$ put disproportionally more density to some regions of the latent space than to others. Thus, to prevent that, we propose to train the latent space so that the distribution of cliques attain broad coverage.
@@
Individual cliques of the learned representations should have broad coverage.
@@
To meet this requirement, we leverage tools from representation learning, but in a novel way. Namely, we put a variational bottleneck (Kingma & Welling, 2013; Higgins et al., 2016; Alemi et al., 2016, VIB) on individual cliques of our representations that brings their distribution closer to a prior with wide coverage, which we choose to be the standard-normal prior. To implement it, when computing the loss for a single example, we sample a single clique to compute the VIB for at random, as opposed to computing it for the joint latent variable like in the classical VIB,
$\text{VIB}({\mathbf{x}},\theta)=\mathbb{E}_{i\sim U[N_{\text{clique}}]}\big{[}\text{KL}\big{(}e_{\theta}({\mathbf{z}}_{C_{i}}|{\mathbf{x}}),p_{C_{i}}({\mathbf{z}}_{C_{i}})\big{)}\big{]},$
@@
where $e_{\theta}({\mathbf{z}}_{C}|{\mathbf{x}})$ is the density of clique $C$ produced by our learnable encoder $e_{\theta}({\mathbf{z}}|{\mathbf{x}})$, and $p_{C}({\mathbf{z}}_{C})$ is the density of ${\mathbf{z}}_{C}$ under the standard-normal distrituion. Note that it is not equivalent to the classical, down-weighted VIB in expectation either since our cliques overlap, meaning that knots contribute to the VIB more often than regular dimensions. We ablate for the impact of this term in Appendix C.
Lastly, since in practical problems, such as biological and chemical design, the data is often discrete (Uehara et al., 2024a; Yang et al., 2024), the designs cannot always be readily optimized with gradient-based methods. Thus, to be generally applicable, our model should be equipped with a data type-agnostic, approximately invertible map from the data space to a continuous space, where gradient-based optimization is feasible. This brings us to our last desideratum.
Desideratum 3.
The model should have a decoder that maps representations back to the original data space.
Thus, together with the model $f_{\theta}({\mathbf{z}})$ and the encoder $e_{\theta}({\mathbf{z}}|{\mathbf{x}})$, we train a decoder $d_{\theta}({\mathbf{x}}|{\mathbf{z}})$ that reconstructs the designs from the latent variables. Putting it all together, the training objective of our model is a VIB-style likelihood objective with a regression term,
@@
$\vspace{-20pt}L_{\text{clique}}(\theta)$ $=\mathbb{E}_{({\mathbf{x}},{\textnormal{y}})\sim\mathcal{D},{\mathbf{z}}\sim e_{\theta}(\cdot|{\mathbf{z}}),i\sim U[N_{\text{clique}}]}\Big{[}\text{VIB}({\mathbf{x}},\theta)$
$-\log d_{\theta}({\mathbf{x}}|{\mathbf{z}})+\tau\cdot\big{(}{\textnormal{y}}-f_{\theta}({\mathbf{z}})\big{)}^{2}\Big{]},$ (4)
where $\tau$ is a positive coefficient that we set to $10$.
@@
Figure 5: Illustration of information flow in Cliqueformer’s training. Data are shown in navy, learnable variables in blue, neural modules in red, and loss functions in pink. The input ${\mathbf{x}}$ is passed to a transformer encoder to compute representation ${\mathbf{z}}$ which is decomposed into cliques with small overlapping knots (highlighted in colors on the figure). The representation goes to the parallel MLPs whose outputs, added together, predict target y. The representation ${\mathbf{z}}$ is also fed to a transformer decoder that tries to recover the original input ${\mathbf{x}}$. Additionally, the representation goes through an information bottleneck during training.
@@
We use transformer networks (Vaswani et al., 2017) for both encoder $e_{\theta}$ and decoder $d_{\theta}$ to achieve the expressivity needed for FGM decomposition in latent space. The encoder transforms input ${\mathbf{x}}\in\mathbb{R}^{d}$ into $d$ vectors of size $d_{\text{model}}$, processes them as token embeddings, and outputs a normal distribution $e_{\theta}({\mathbf{z}}|{\mathbf{x}})$. Sampled representations ${\mathbf{z}}$ are arranged into $N_{\text{clique}}$ cliques (dimension $d_{\text{clique}}$, knot size $d_{\text{knot}}$) before being passed to both the predictive model $f_{\theta}$ and transformer decoder $d_{\theta}$. For illustration of the information flow in Cliqueformer’s training consult Figure 5.
@@
3:     Take a gradient step on the parameter $\theta$ with respect to $L_{\text{clique}}(\theta)$ from Equation (3.2).
@@
5:  Sample $B$ examples ${\mathbf{x}}^{(i_{b})}\sim\mathcal{D}$ from the dataset.
6:  Encode the examples ${\mathbf{z}}^{(i_{b})}\sim e_{\theta}({\mathbf{z}}|{\mathbf{x}}^{(i_{b})})$,.
@@
8:     Decay the representations, ${\mathbf{z}}^{(i_{b})}\leftarrow(1-\lambda){\mathbf{z}}^{(i_{b})}$.
@@ 3.3 Optimizing Designs With Cliqueformer
Once Cliqueformer is trained, we use it to optimize new designs. Typically, MBO methods initialize this step at a sample of designs $({\mathbf{x}}^{i_{b}})_{b=1}^{B}$ drawn from the dataset (Trabucco et al., 2021). Since in our algorithm the optimization takes place in the latent space $\mathcal{Z}$, we perform this step by encoding the sample of designs with Cliqueformer’s encoder, ${\mathbf{z}}^{i_{b}}\sim e_{\theta}({\mathbf{z}}|{\mathbf{x}}^{i_{b}})$. We then optimize the representation ${\mathbf{z}}^{i_{b}}$ of design ${\mathbf{x}}^{i_{b}}$ to maximize our model’s value,
$\vspace{-10pt}L_{\text{mbo}}\big{(}({\mathbf{z}}^{i_{b}})_{b=1}^{B}\big{)}=\frac{1}{B}\sum_{b=1}^{B}f_{\theta}({\mathbf{z}}^{i_{b}}),\vspace{-10pt}$ (5)
at the same time minding the enumerator of the regret bound from Theorem 1. That is, we don’t want the optimizer to explore regions under which the marginal densities $e_{\theta}({\mathbf{z}}_{C})=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}[e_{\theta}({\mathbf{z}}_{C}|{\mathbf{x}})]$ are small. Fortunately, since the encoder was trained with standard-normal prior on the cliques, $p({\mathbf{z}}_{C})=N(0_{d_{\text{clique}}},I_{d_{\text{clique}}})$, we know that values of ${\mathbf{z}}$ closer to the origin have unilaterally higher marginals. This simple property of standard-normal distribution allows us to confine the optimizer’s exploration to designs with in-distribution cliques by exponentially decaying the design at every optimization step. Thus, we use AdamW as our optimizer (Loshchilov et al., 2017) (see Appendix C for an ablation study). We provide the pseudocode of the whole procedure of designign with Cliqueformer in Algorithm 1.
@@ 4 Related Work
The idea of using machine learning models in optimization problems has existed for a long time, and has been mainly cultivated in the literature on Bayesian optimization (Williams & Rasmussen, 2006; Brochu et al., 2010; Snoek et al., 2012, BO). The BO paradigm relies on two core assumptions: availability of data of examples paired with their target function values, as well as access to an oracle that allows a learning algorithm to query values of proposed examples. Thus, similarly to reinforcement learning, the challenge of BO is to balance exploitation and exploration of the black-box function modeled by a Gaussian process.
@@
Recently, to help BO tackle very high-dimensional problems, techniques of decomposing the target function have become more popular (Kandasamy et al., 2015; Rolland et al., 2018). Most commonly, these methods decompose the target function into functions defined on the input’s partitions. While such models are likely to deviate far from the functions’ ground-truth structure, one can derive theoretical guarantees for a range of such decompositions under the BO’s query budget assumptions (Ziomek & Bou-Ammar, 2023). However, these results do not apply to our setting of offline MBO, where no additional queries are available and the immediate reliability on the model is essential. Furthermore, instead of partitioning the input, we decompose our prediction over a latent variable that is learned by a transformer, enabling the model to acquire an expressive structure over which the decomposition is valid.
@@
Offline model-based optimization (MBO) has gained traction in domains where BO assumptions cannot be met, offering solution generation from static datasets without additional queries. Early applications in molecule design (Gómez-Bombarelli et al., 2018) utilized variational auto-encoders (Kingma & Welling, 2013, VAE) to learn continuous molecular representations. While this demonstrated deep learning’s potential, it didn’t explore the target function’s structural properties. The Conditioning by Adaptive Sampling (CBaS) algorithm (Brookes et al., 2019) introduced a data type-agnostic approach for refining designs using non-differentiable oracle predictions. While this can be combined with trainable models through auto-focusing (Fannjiang & Listgarten, 2020), our setting differs by assuming a differentiable neural network model of the black-box function, enabling optimization through automatic differentiation (Paszke et al., 2019). Trabucco et al. (2021) introduced a neural network-based method, exactly for our setting, dubbed Conservative Objective Models (COMs), where a surrogate model is trained to both predict values of examples that can be found in the dataset, and penalize those that are not. COMs differs from our work fundamentally, since its contribution lies in the formulation of the conservative regularizer applied to arbitrary neural networks, while we focus on scalable model architectures that facilitate computational design. Another recent line of work proposes to tackle the design problem through means of generative modeling. BONET (Mashkaria et al., 2023) and DDOM (Krishnamoorthy et al., 2023) are examples of works that bring the most recent novelties of the field to address design tasks. BONET does so by training a transformer to generate sequences of designs that monotonically improve in their value, and DDOM by training a value-conditioned diffusion model. That is, these methods attempt to generate high-value designs through novel conditional generation mechanisms. Instead, we model MBO as a maximization problem, and propose a scalable model that acquires the structure of the black-box function through standard gradient-based learning. To this end, we bring powerful techniques from deep learning and generative modeling, like transformers (Vaswani et al., 2017) and variational-information bottlenecks (Kingma & Welling, 2013; Alemi et al., 2016).
Task Grad.Asc. RWR COMs DDOM Transformer Cliqueformer
Lat. RBF 11 0.00 $\pm$ 0.00 (0%) 0.08 $\pm$ 0.08 (1%) 0.66 $\pm$ 0.04 (72%) 0.00 $\pm$ 0.00 (0%) 0.47 $\pm$ 0.05 (60%) 0.65 $\pm$ 0.07 (74%)
Lat. RBF 31 0.00 $\pm$ 0.00 (0%) 0.31 $\pm$ 0.10 (3%) 0.50 $\pm$ 0.05 (32%) 0.00 $\pm$ 0.00 (0%) 0.00 $\pm$ 0.00 (0%) 0.64 $\pm$ 0.05 (76%)
Lat. RBF 41 0.00 $\pm$ 0.00 (0%) 0.35 $\pm$ 0.08 (3%) 0.45 $\pm$ 0.06 (16%) 0.00 $\pm$ 0.00 (0%) 0.20 $\pm$ 0.01 (75%) 0.66 $\pm$ 0.05 (72%)
Lat. RBF 61 0.00 $\pm$ 0.00 (0%) 0.29 $\pm$ 0.10 (4%) 0.25 $\pm$ 0.04 (7%) 0.00 $\pm$ 0.00 (0%) 0.16 $\pm$ 0.03 (64%) 0.66 $\pm$ 0.05 (68%)
Superconductor 1.13 $\pm$ 0.08 1.03 $\pm$ 0.07 0.97 $\pm$ 0.08 1.22 $\pm$ 0.08 0.96 $\pm$ 0.05 1.43 $\pm$ 0.04
TF-Bind-8 0.99 $\pm$ 0.00 1.58 $\pm$ 0.03 1.57 $\pm$ 0.02 1.55 $\pm$ 0.03 1.48 $\pm$ 0.03 1.58 $\pm$ 0.01
DNA HEPG2 2.16 $\pm$ 0.07 1.91 $\pm$ 0.12 1.20 $\pm$ 0.09 1.82 $\pm$ 0.10 2.13 $\pm$ 0.06 2.10 $\pm$ 0.07
DNA k562 2.11 $\pm$ 0.06 1.91 $\pm$ 0.11 1.80 $\pm$ 0.12 2.61 $\pm$ 0.21 2.60 $\pm$ 0.19 3.15 $\pm$ 0.07
Ave.score $\uparrow$ 0.80 0.93 0.93 0.90 1.00 1.36
Ave.rank $\downarrow$ 4.75 3.88 3.63 4.50 4.25 1.38
Table 1: Experimental results of Cliqueformer and the baselines. Each score is the mean of values of top_k=10 of batch_size=1000 designs, averaged over 5 runs. The values were normalized with the min-max scheme, where the minimum and the maximum are taken from the dataset, so that the scores of the designs in the dataset are in range $[0,1]$. We note that, unlike Trabucco et al. (2022), we take the maximum from the data available for the MBO model training (the union of the train and test data), and not from the oracle training data, to make the results more interpretable (see Appendix D). We also provide the standard deviation estimates as the standard deviation for each of the top_k samples, averaged over the runs. Additionally, for Lat. RBF tasks, we provide the average validity in parentheses, calculated as the percentage of valid designs from the batch_size produced candidates, averaged across the runs. We provide average score (the higher the better) and the average rank (the lower the better) for each method.
@@
The work of Grudzien et al. (2024) introduced the theoretical foundations of functional graphical models (FGMs), including Theorem 1. However, as we have shown in Theorem 2, their graph discovery heuristic for neural networks renders learning the black-box function’s structure an open problem. We addressed this issue by subsuming the graph discovery step in the architecture of our Cliqueformer that learns to abide by a pre-defined FGM. As such, the model learns the structure of the target function, as well as learns to predict its value, in synergy within an end-to-end training. For more related works, please refer to Appendix A.